OpenStudy (anonymous):

What is the slope of the line y = 3?? how do i find this answer/??

jimthompson5910 (jim_thompson5910):

hint: y = 3 is the same as y = 0x + 3

jimthompson5910 (jim_thompson5910):

compare y = 0x + 3 to y = mx+b

OpenStudy (anonymous):

so it's just 0 right?? or is it undefined???

OpenStudy (austinl):

Slope is defined as \[m = \frac{\Delta y}{\Delta x}\rightarrow \frac{rise}{run}\] For every change in x, is there any change in y?

OpenStudy (jdoe0001):

for a horizontal line the slope is 0 \(\bf slope = \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\implies\cfrac{3 -3}{\textit{some value here}} \implies \cfrac{0}{\textit{some value here}}\) anything dividing 0 = 0 0/25 = 0 0/1000000 = 0 0/whatever = 0